4x - 9/3 + x/6 = x/2 - 2
Answer by stanbon(75887)

4x - 9/3 + x/6 = x/2 - 2
-------
Multiply thru by 6 to get:
24x - 18 + x = 3x - 12
-------
22x = 6
x = 3/11
----------------------------
Cheers,
Stan H.
====================